Evaluate the integral of (3x² - 4 +1) dx


1
Expert's answer
2021-03-24T13:42:09-0400

We have to find integral "\int" (3x2 - 4 + 1) dx

( since we know that "\int" xn dx = "\frac{x^n +1}{n+1}" )

therefore,

"\int" (3x2 - 4 + 1) dx

= "\int" (3x2 - 3) dx

= "\int" 3x2 dx + "\int"-3 dx

= 3"\int"x2 dx - 3"\int"1 dx

= 3"\frac{x^3}{3}" - 3x + c

= x3 - 3x + c

where c is any constant